1. On the moon, an eagle feather dropped by Apollo astronauts falls with an acceleration of 1.6 m.s^-2. Ted, the astronaut who has done this experiment, has a mass of 79 kg on Earth. Using these values answer the following questions:

a) What is Ted's weight (on Earth)?
b) What is his weight on the moon?
c) What is his mass on the moon?

2. My bathroom scales read in kilograms, meaning that, if I put a mass m on top of them, in mechanical equilibrium, the dial (accurately) shows the mass m in kg. The scales are light.
I apply a force of 85 newtons to the scales. What do they read?
1
Expert's answer
2020-03-05T09:40:52-0500

1) a)


"W_E=mg_E=(79)(9.8)=770\ N"

b)


"W_M=mg_M=(79)(1.6)=130\ N"

c)


"m_M=m_E=m=79\ kg"

2)


"m=\frac{W}{g}=\frac{85}{9.8}=8.7\ kg"